# Well-posedness and stability of the Lagrange representation of the n-D wave equation via boundary triples

### v1

@@ 2.1. Sobolev Spaces
Note that $\mathrm{C}_{\mathrm{c}}^{\infty}(\overline{\Omega})\coloneqq\left\{f\big{|}_{\overline{\Omega}}\,\middle|\,\mathopen{}f\in\mathrm{C}_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})\right\}$ is dense in $\mathrm{H}(\operatorname{div},\Omega)$, see, e.g., [7, Ch. IX Part A Sec. 2 Thm. 1] or [20, Thm. 3.18].
2.2. Trace Operators
The following can be found in more detail in [15, Appendix A]. For $f\in\mathrm{C}_{\mathrm{c}}^{\infty}(\overline{\Omega})$ we can define the map
By continuous extension we can extend $\tilde{\gamma}_{0}$ to a continuous map $\gamma_{0}\colon\mathrm{H}^{1}(\Omega)\to\mathrm{L}^{2}(\partial\Omega)$, the Dirichlet trace operator. We define its image as
$\mathrm{H}^{\frac{1}{2}}(\partial\Omega)\coloneqq\operatorname{ran}\gamma_{0}\quad\text{equipped with }\quad\lVert\phi\rVert_{\mathrm{H}^{\frac{1}{2}}(\partial\Omega)}\coloneqq\inf\{\lVert g\rVert_{\mathrm{H}^{1}(\Omega)}\,|\,\mathopen{}\gamma_{0}g=\phi\}.$
This space is even a Hilbert space as it can be represented by the quotient space $\mathrm{H}^{1}(\Omega)/\ker\gamma_{0}$. Note that $\gamma_{0}$ is a continuous map from $\mathrm{H}^{1}(\Omega)$ to $\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$ and $\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$ is continuously embedded in $\mathrm{L}^{2}(\partial\Omega)$. Let $\Gamma_{0},\tilde{\Gamma}\subseteq\partial\Omega$ be a (measurable) partition of $\partial\Omega$ (up to sets of measure zero). We can restrict $\gamma_{0}f$ to $\Gamma_{0}$, which gives that $\gamma_{0}\big{|}_{\Gamma_{0}}\colon f\mapsto\gamma_{0}f\big{|}_{\Gamma_{0}}$ is continuous from $\mathrm{H}^{1}(\Omega)$ to $\mathrm{L}^{2}(\Gamma_{0})$ also. Hence,
$\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)\coloneqq\ker\gamma_{0}\big{|}_{\Gamma_{0}}=\{f\in\mathrm{H}^{1}(\Omega)\,|\,\mathopen{}\gamma_{0}f\big{|}_{\Gamma_{0}}=0\}$
is closed and therefore a Hilbert space equipped with $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{H}^{1}(\Omega)}$. We define the corresponding trace space
$\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})\coloneqq\gamma_{0}\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\partial\Omega)$
and endow it with $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{H}^{\frac{1}{2}}(\partial\Omega)}$. As the functions in $\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})$ are 00 on $\Gamma_{0}$ we usually regard this space as continuously embedded in $\mathrm{L}^{2}(\tilde{\Gamma})$. We define the dual space of $\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})$ with pivot space $\mathrm{L}^{2}(\tilde{\Gamma})$ as
$\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})\coloneqq\bigl{(}\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})\bigr{)}^{\prime},$
i.e., $\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})\subseteq\mathrm{L}^{2}(\tilde{\Gamma})\subseteq\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})$ forms a Gelfand triple. The pairing between $\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})$ and $\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})$ is an extension of the inner product of $\mathrm{L}^{2}(\tilde{\Gamma})$:
$\langle f,g\rangle_{\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma}),\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})}\coloneqq\lim_{n\to\infty}\langle f_{n},g\rangle_{\mathrm{L}^{2}(\tilde{\Gamma})},$
where $(f_{n})_{n\in\mathbb{N}}$ is a sequence in $\mathrm{L}^{2}(\tilde{\Gamma})$ converging to $f\in\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})$ (w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})}$) and $g\in\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})$. Moreover, we define $\langle g,f\rangle_{\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma}),\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})}$ as the complex conjugate of $\langle f,g\rangle_{\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma}),\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})}$ and we will use the short notation
$\langle f,g\rangle_{\mp\frac{1}{2}}\coloneqq\langle f,g\rangle_{\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma}),\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})}\quad\text{and}\quad\langle g,f\rangle_{\pm\frac{1}{2}}\coloneqq\langle g,f\rangle_{\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma}),\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})}.$
For $f\in\mathrm{C}_{\mathrm{c}}^{\infty}(\overline{\Omega})^{d}$ we can define the map
$\tilde{\gamma}_{\nu}\colon\left\{\begin{array}[]{rcl}\mathrm{C}_{\mathrm{c}}^{\infty}(\overline{\Omega})^{d}&\to&\mathrm{L}^{2}(\tilde{\Gamma}),\\ f&\mapsto&\nu\cdot f\big{|}_{\tilde{\Gamma}}.\end{array}\right.$
By continuous extension we can extend $\tilde{\gamma}_{\nu}$ to $\gamma_{\nu}\colon\mathrm{H}(\operatorname{div},\Omega)\to\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})$, the normal trace operator.
Theorem 2.1 ([15, Theorem A.8]).
The normal trace operator $\gamma_{\nu}\colon\mathrm{H}(\operatorname{div},\Omega)\to\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})$ is a bounded, linear and surjective operator.
@@ 2.2. Trace Operators
$\langle\operatorname{div}F,g\rangle_{\mathrm{L}^{2}(\Omega)}=\langle F,\nabla g\rangle_{\mathrm{L}^{2}(\Omega)^{d}}+\langle\gamma_{\nu}F,\gamma_{0}g\rangle_{\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma}),\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})}$
@@
For $F\in\mathrm{C}^{\infty}(\overline{\Omega})^{d}$ we have for $g\in\mathring{\mathrm{H}}_{\Gamma_{0}}^{1}(\Omega)$:
@@
Density of $\mathrm{C}_{\mathrm{c}}^{\infty}(\overline{\Omega})^{d}$ in $\mathrm{H}(\operatorname{div},\Omega)$ implies the claim. ∎
The Friedrichs/Poincaré inequality (Theorem C.1) allows $\mathrm{H}^{1}(\Omega)$ to be equipped with an equivalent inner product:
$\langle f,g\rangle_{\mathrm{H}^{1}(\Omega)}\coloneqq\langle\nabla f,\nabla g\rangle_{\mathrm{L}^{2}(\Omega)}+\langle k_{1}\gamma_{0}f,\gamma_{0}g\rangle_{\mathrm{L}^{2}(\tilde{\Gamma})}.$ (5)
In particular we are interested in $\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)$, which is a subset of $\mathrm{H}^{1}(\Omega)$. Note that (5) gives still an equivalent norm on $\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)$ if $k_{1}=0$, under the condition that $\Gamma_{0}\neq\emptyset$. Hence, we either need
@@ 2.3. Dissipative operators
Let $A$ be a densely defined linear operator on a Hilbert space $H$. The following are equivalent:
@@
    $A$ is dissipative and $\operatorname{ran}(\lambda I-A)$ is onto for some (and hence all) $\lambda>0$.
@@
    $A$ is closed and both $A$ and $A^{\ast}$ are dissipative.
Proofs of these statements as well as the next theorem are contained in [9, Sec. II.3b]. We will rely on the following well-known theorem:
@@
Let $A\colon H\supseteq\operatorname{dom}A\to H$ be a densely defined linear operator on a Hilbert space $H$. Then $A$ is infinitesimal generator of a contraction semigroup if and only if $A$ is maximally dissipative.
@@ 2.4. Helmholtz decomposition
The classical Helmholtz decomposition allows the splitting of a vector field into a gradient field and a divergence-free vector field. We will also use a slight modification of the classical Helmholtz decomposition and introduce
$\mathring{\mathrm{H}}^{1}(\Omega)$ $\coloneqq\mathring{\mathrm{H}}_{\partial\Omega}^{1}(\Omega)=\{f\in\mathrm{H}^{1}(\Omega)\,|\,\mathopen{}\gamma_{0}f=0\}$
$\mathrm{H}(\operatorname{div}0,\Omega)$$\coloneqq\{v\in\mathrm{H}(\operatorname{div},\Omega)\,|\,\mathopen{}\operatorname{div}v=0\}=\ker\operatorname{div}.$
With this preparation we can show the following modification of the classical result:
@@
Let $T\in\mathcal{L}_{\mathrm{b}}(\mathrm{L}^{2}(\Omega))$ be boundedly invertible. Then
@@
where $\oplus_{T^{-1}}$ denotes the orthogonal sum w.r.t. the inner product $\langle f,g\rangle_{T^{-1}}\coloneqq\langle T^{-1}f,g\rangle$.
@@
The standard Helmholtz decomposition with $T=\mathrm{id}$ is simply an application of the orthogonal decomposition of $\mathrm{L}^{2}(\Omega)$ into $\overline{\operatorname{ran}L}$ and $\ker L^{\ast}$, where $L$ is the densely defined linear operator $\mathring{\nabla}$, i.e., the gradient with Dirichlet boundary conditions ($\nabla$ restricted to $\mathring{\mathrm{H}}^{1}(\Omega)\subseteq\mathrm{L}^{2}(\Omega)$).
Hence, we only have to verify closedness of $\operatorname{ran}L$. Let $\{u_{n}\}_{n}\subseteq\operatorname{ran}L$ with $u_{n}=Lv_{n}$ for some $v_{n}\in\mathrm{L}^{2}(\Omega)$ converge to some $u\in\mathrm{L}^{2}(\Omega)$. Convergence of $(u_{n})_{n}$ implies boundedness of $(Lv_{n})_{n}$ in $\mathrm{L}^{2}(\Omega)$. Since $\lVert\nabla\boldsymbol{\cdot}\rVert_{\mathrm{L}^{2}(\Omega)}$ and $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{H}^{1}(\Omega)}$ are equivalent on $\mathring{\mathrm{H}}^{1}(\Omega)$ by virtue of the standard Poincaré inequality, this implies boundedness of $(v_{n})_{n}$ as a sequence in $\mathrm{H}^{1}(\Omega)$. By Rellich’s theorem, there exists a strongly convergent subsequence of $(v_{n})_{n}$ in $\mathrm{L}^{2}(\Omega)$. Appealing to the closedness of $L$, we obtain that $u=\lim_{n\to\infty}u_{n}=\lim_{n\to\infty}Lv_{n}=L\lim_{n\to\infty}v_{n}$, hence $u\in\operatorname{ran}L$.
The $T$ in front of the first orthogonal space and the $T^{-1}$ attached in the first component of the inner product corresponding to $\oplus_{T^{-1}}$ cancel out and give the standard Helmholtz decomposition. ∎
@@ 2.5. Boundary triples
Boundary triples were originally developed for symmetric operators see [12, 4], but they can be equivalently defined in the skew-symmetric case, see [23] and [19, Ch. 2.4]. This is a simple consequence of multiplying the equations by the imaginary unit $\mathrm{i}$. We will present the version from [19] that allows dual pairs as the boundary spaces instead of a single (identified) Hilbert space. There is also the notion of m-boundary tuples [8], which additionally requires a pivot space for the dual pair.
@@
    $B=\begin{pmatrix}B_{1}\\ B_{2}\end{pmatrix}\colon\left\{\begin{array}[]{rcl}\operatorname{dom}(A_{0}^{\ast})&\to&\mathcal{B}_{+}\times\mathcal{B}_{-},\\ x&\mapsto&\begin{pmatrix}B_{1}x\\ B_{2}x\end{pmatrix}\end{array}\right.$
    is onto.
@@
    The following abstract Green’s identity holds for all $x,y\in\operatorname{dom}(A_{0}^{\ast})$:
@@
The abstract Green’s identity (ii) formalizes an integration by parts formula, which is the reason for the designation boundary triple.
A boundary triple enables us to parameterize all boundary conditions such that $A_{0}^{\ast}$ restricted to all elements of $\operatorname{dom}A_{0}^{\ast}$, that satisfy the boundary condition, is maximally dissipative. In our situation we will obtain a boundary triple that involves the trace spaces $\mathring{\mathrm{H}}^{\frac{1}{2}}(\Gamma)$ and $\mathrm{H}^{-\frac{1}{2}}(\Gamma)$, but we would rather formulate boundary conditions in the pivot space $\mathrm{L}^{2}(\Gamma)$. Verifying dissipativity is usually straightforward, but the maximality can be tricky. In the next section, we will arrive at a situation in which $((\mathcal{B}_{-},\mathcal{B_{+}}),B_{1},B_{2})$ is a boundary triple [^8] The order of $(\mathcal{B}_{-},\mathcal{B_{+}})$ is swapped on purpose, because later we will use $(\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma}),\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma}))$., where the duality of $(\mathcal{B}_{+},\mathcal{B_{-}})$ is induced by a pivot space $\mathcal{B}_{0}$. Moreover, $\mathcal{B_{+}}\subseteq\mathcal{B}_{0}\subseteq\mathcal{B}_{-}$ forms a Gelfand triple. We want to formulate Robin type boundary conditions of the form
@@
where $\Theta\in\mathcal{L}_{\mathrm{b}}(\mathcal{B}_{0})$. So strictly speaking we have to take the embedding mappings $j_{+}\colon\mathcal{B}_{+}\to\mathcal{B}_{0}$ and $j_{-}\colon\mathcal{B}_{0}\to\mathcal{B}_{-}$ into account, i.e.,
$j_{-}^{-1}B_{1}x+\Theta j_{+}B_{2}x=0$
@@
[…]
@@
$\hat{\Theta}\coloneqq-j_{-}\Theta j_{+}.$
By the theory of boundary triples $A_{\Theta}$ is maximally dissipative if and only if $\hat{\Theta}$ is maximally dissipative, see, e.g., [19, Prop. 2.4.10] or [4, Cor. 2.1.4.]. [^9] In the second reference boundary triples are introduced for the symmetric setting.
@@
Let $(\mathcal{B}_{+},\mathcal{B}_{0},\mathcal{B}_{-})$ be a Gelfand triple, $((\mathcal{B}_{-},\mathcal{B_{+}}),B_{1},B_{2})$ be a boundary triple for $A_{0}^{\ast}$ and $\Theta\in\mathcal{L}_{\mathrm{b}}(\mathcal{B}_{0})$. Then $A_{\Theta}$ is maximally dissipative, if $\Theta$ is positive (semi-definite), i.e., $\langle\Theta h,h\rangle_{\mathcal{B}_{0}}\geq 0$ for all $h\in\mathcal{B}_{0}$.
@@
Note that $A_{\Theta}$ is maximally dissipative, if $\hat{\Theta}\coloneqq-j_{-}\Theta j_{+}$ is maximally dissipative. Hence, it is sufficient to show (the stronger assertion) that $\hat{\Theta}$ is self-adjoint and negative.
@@
$\hat{\Theta}^{\ast}=-(j_{-}\Theta j_{+})^{\ast}=-j_{+}^{\ast}\Theta^{\ast}j_{-}^{\ast}=-(j_{-}\Theta j_{+})^{\ast}=\hat{\Theta},$
which implies that $\hat{\Theta}$ is self-adjoint and therefore maximally dissipative. Hence, also $\hat{\Theta}^{-1}$ is maximally dissipative (in the sense of linear relations). Finally, we can apply [19, Prop. 2.4.10], which deduces the maximal dissipativity of $A_{\Theta}$ from the maximal dissipativity of $\hat{\Theta}^{-1}$ (Note that [19] uses a different notation. The operator $A_{\Theta}$ corresponds to $A_{\hat{\Theta}^{-1}}$ in the notation of [19]). ∎

### v2

@@ 2.2. Trace Operators
$\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})\coloneq\gamma_{0}\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\partial\Omega)$
[…]
$\langle f,g\rangle_{\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma}),\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})}\coloneq\lim_{n\to\infty}\langle f_{n},g\rangle_{\mathrm{L}^{2}(\tilde{\Gamma})}$,
[…]
$\langle f,g\rangle_{\mp\frac{1}{2}}\coloneq\langle f,g\rangle_{\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma}),\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})}\quad\text{and}\quad\langle g,f\rangle_{\pm\frac{1}{2}}\coloneq\langle g,f\rangle_{\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma}),\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma})}$.
For $f\in\mathrm{C}_{\mathrm{c}}^{\infty}(\overline{\Omega};\mathbb{C}^{d})$ we can define the map
[…]
Theorem 2.1 ([18, Theorem A.8]).
[…]
@@
$\langle\operatorname{div}F,g\rangle_{\mathrm{L}^{2}(\Omega)}+\langle F,\nabla g\rangle_{\mathrm{L}^{2}(\Omega)}=\langle\gamma_{\nu}F,\gamma_{0}g\rangle_{\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma}),\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma})}$.
@@
For $F\in\mathrm{C}^{\infty}(\overline{\Omega})^{d}$ and $g\in\mathring{\mathrm{H}}_{\Gamma_{0}}^{1}(\Omega)$ we have:
@@
Density of $\mathrm{C}_{\mathrm{c}}^{\infty}(\overline{\Omega};\mathbb{C}^{d})$ in $\mathrm{H}(\operatorname{div},\Omega)$ implies the claim. ∎
The Friedrichs/Poincaré inequality (Theorem A.1) allows $\mathrm{H}^{1}(\Omega)$ to be equipped with an equivalent inner product:
$\langle f,g\rangle_{\mathrm{H}^{1}(\Omega)}\coloneq\langle\nabla f,\nabla g\rangle_{\mathrm{L}^{2}(\Omega)}+\langle k_{1}\gamma_{0}f,\gamma_{0}g\rangle_{\mathrm{L}^{2}(\tilde{\Gamma})}$. (5)
In particular we are interested in $\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)$, which is a subset of $\mathrm{H}^{1}(\Omega)$. Note that (5) still defines an equivalent norm on $\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)$ if $k_{1}=0$, provided that $\Gamma_{0}\neq\emptyset$. Hence, we need
@@ 2.3. Dissipative operators
Let $A$ be a densely defined, linear operator on a Hilbert space $H$. The following are equivalent:
@@
    AA is dissipative and $\operatorname{ran}(\lambda I-A)$ is onto for some (and hence all) $\lambda>0$.
@@
    AA is closed and both $A$ and $A^{\ast}$ are dissipative.
Proofs of these statements as well as the next theorem are contained in [12, Sec. II.3b]. We will rely on the following well-known theorem:
@@
Let $A\colon H\supseteq\operatorname{dom}A\to H$ be a densely defined, linear operator on a Hilbert space $H$. Then $A$ is infinitesimal generator of a contraction semigroup if and only if $A$ is maximally dissipative.
@@ 2.4. Helmholtz decomposition
The classical Helmholtz decomposition allows the splitting of a vector field into a gradient field and a divergence-free vector field. We will use a slight modification of the classical Helmholtz decomposition and introduce
$\mathring{\mathrm{H}}^{1}(\Omega)$ $\coloneq\mathring{\mathrm{H}}_{\partial\Omega}^{1}(\Omega)=\{f\in\mathrm{H}^{1}(\Omega)\,|\,\mathopen{}\gamma_{0}f=0\}$
$\mathrm{H}(\operatorname{div}0,\Omega)$ $\coloneq\{v\in\mathrm{H}(\operatorname{div},\Omega)\,|\,\mathopen{}\operatorname{div}v=0\}=\ker\operatorname{div}.$
With this notation, we show the following modification of the classical result:
@@
Let $T\in\mathcal{L}_{\mathrm{b}}(\mathrm{L}^{2}(\Omega))$ be as in (A2), i.e., $T$ positive definite such that $c^{-1}\mathrm{I}<T<c\mathrm{I}$ for some $c>0$. Then
@@
where $\oplus_{T^{-1}}$ denotes the orthogonal sum w.r.t. the inner product $\langle f,g\rangle_{T^{-1}}\coloneq\langle T^{-1}f,g\rangle$.
@@
[…]
Hence, we only have to verify closedness of $\operatorname{ran}L$. Let $\{u_{n}\}_{n}\subseteq\operatorname{ran}L$ with $u_{n}=Lv_{n}$ for some $v_{n}\in\mathrm{L}^{2}(\Omega)$ converge to some $u\in\mathrm{L}^{2}(\Omega)$. Convergence of $(u_{n})_{n}$ implies boundedness of $(Lv_{n})_{n}$ in $\mathrm{L}^{2}(\Omega)$. Since $\lVert\nabla\boldsymbol{\cdot}\rVert_{\mathrm{L}^{2}(\Omega)}$ and $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{H}^{1}(\Omega)}$ are equivalent on $\mathring{\mathrm{H}}^{1}(\Omega)$ by virtue of the standard Friedrichs/Poincaré inequality, this implies boundedness of $(v_{n})_{n}$ as a sequence in $\mathrm{H}^{1}(\Omega)$. By the Rellich–Kondrachov theorem ([19, Thm. 12.30]), there exists a strongly in $\mathrm{L}^{2}(\Omega)$ convergent subsequence $(v_{n_{k}})_{k}$ of $(v_{n})_{n}$. Appealing to the closedness of $L$, we obtain that $u=\lim_{k\to\infty}u_{n_{k}}=\lim_{k\to\infty}Lv_{n_{k}}=L\lim_{k\to\infty}v_{n_{k}}$, hence $u\in\operatorname{ran}L$. ∎
@@ 2.5. Boundary triples
Boundary triples were originally developed for symmetric operators, cf. [15, 6], but they can be equivalently defined in the skew-symmetric case, cf. [29] and [24, Ch. 2.4]. This is a simple consequence of multiplying the equations by the imaginary unit $\mathrm{i}$.
Boundary triples are certainly not the only tool one can apply to model (dissipative) boundary conditions. Other viable concepts include m-boundary tuples (cf. [11]), which additionally require a pivot space for the dual pair, the recent notion of generalized boundary triples (cf. [3]), as well as boundary quadruples (cf. [2]) and boundary systems (cf. [22]), which exist even if a skew-adjoint extension of the operator does not exist. However, all these notions are equivalent to the concept of boundary triples given strong enough assumptions (cf. [28] for the equivalence of boundary systems and boundary triples). In our context, it suffices to consider boundary triples though. We will present the definition used in [24], which allows a dual pair instead of a single (identified) Hilbert space as a boundary space.
@@
[…]
    is onto and
@@
    the following abstract Green identity holds for all $x,y\in\operatorname{dom}(A_{0}^{\ast})$:
@@
The abstract Green identity (ii) formalizes an integration by parts formula, which is the reason for the designation boundary triple.
A boundary triple enables us to parameterize all boundary conditions such that $A_{0}^{\ast}$ restricted to all elements of $\operatorname{dom}A_{0}^{\ast}$, that satisfy the boundary condition, is maximally dissipative. In our situation we will obtain a boundary triple that involves the trace spaces $\mathring{\mathrm{H}}^{\frac{1}{2}}(\Gamma)$ and $\mathrm{H}^{-\frac{1}{2}}(\Gamma)$, but we would rather formulate boundary conditions in the pivot space $\mathrm{L}^{2}(\Gamma)$. Verifying dissipativity is usually straightforward, but the maximality can be tricky. In the next section, we will arrive at a situation in which $((\mathcal{B}_{-},\mathcal{B_{+}}),B_{1},B_{2})$ is a boundary triple [^9] The order of $(\mathcal{B}_{-},\mathcal{B_{+}})$ is swapped on purpose, because later we will use $)(\mathrm{H}^{-\frac{1}{2}}(\tilde{\Gamma}),\mathring{\mathrm{H}}^{\frac{1}{2}}(\tilde{\Gamma}$))., where the duality of $(\mathcal{B}_{+},\mathcal{B_{-}})$ is induced by a pivot space $\mathcal{B}_{0}$. Moreover, $\mathcal{B_{+}}\subseteq\mathcal{B}_{0}\subseteq\mathcal{B}_{-}$ forms a Gelfand triple. We want to formulate Robin type boundary conditions of the form
@@
where $\Theta\in\mathcal{L}_{\mathrm{b}}(\mathcal{B}_{0})$. To formulate this rigorously, we have to take the embedding mappings $j_{+}\colon\mathcal{B}_{+}\to\mathcal{B}_{0}$ and $j_{-}\colon\mathcal{B}_{0}\to\mathcal{B}_{-}$ into account, i.e.,
$j_{-}^{-1}B_{1}x+\Theta j_{+}B_{2}x=0,$
@@
$\begin{aligned} A_{\Theta}&\coloneq A_{0}^{\ast}\big|_{\operatorname{dom}A_{\Theta}},\\ \operatorname{dom}A_{\Theta}&\coloneq\{x\in\operatorname{dom}A_{0}^{\ast}\,|\,\mathopen{}B_{1}x=-j_{-}\Theta j_{+}B_{2}x\}.\end{aligned}$ (6)
@@
$\hat{\Theta}\coloneq-j_{-}\Theta j_{+}.$
By the theory of boundary triples, $A_{\Theta}$ is maximally dissipative if and only if $\hat{\Theta}$ is maximally dissipative, see, e.g., [24, Prop. 2.4.10] or [6, Cor. 2.1.4].
@@
Let $(\mathcal{B}_{+},\mathcal{B}_{0},\mathcal{B}_{-})$ be a Gelfand triple, $((\mathcal{B}_{-},\mathcal{B_{+}}),B_{1},B_{2})$ be a boundary triple for $A_{0}^{\ast}$ and $\Theta\in\mathcal{L}_{\mathrm{b}}(\mathcal{B}_{0})$. Then $A_{\Theta}$ is maximally dissipative if $\Theta$ is positive (semi-definite), i.e., $\langle\Theta h,h\rangle_{\mathcal{B}_{0}}\geq 0$ for all $h\in\mathcal{B}_{0}$.
Note that instead of positivity we could have asked for accretivity, i.e., $\operatorname{Re}\langle\Theta h,h\rangle_{\mathcal{B}_{0}}\geq 0$ for all $h\in\mathcal{B}_{0}$.
@@
Note that $A_{\Theta}$ is maximally dissipative if $\hat{\Theta}\coloneq-j_{-}\Theta j_{+}$ is maximally dissipative. Hence, it is sufficient to show (the stronger assertion) that $\hat{\Theta}$ is self-adjoint and negative.
@@
$\hat{\Theta}^{\ast}=-(j_{-}\Theta j_{+})^{\ast}=-j_{+}^{\ast}\Theta^{\ast}j_{-}^{\ast}=-j_{-}\Theta j_{+}=\hat{\Theta}$,
which implies that $\hat{\Theta}$ is self-adjoint and therefore maximally dissipative. Hence, $\hat{\Theta}^{-1}$ is maximally dissipative (in the sense of linear relations). Finally, we can apply [24, Prop. 2.4.10], which deduces the maximal dissipativity of $A_{\Theta}$ from the maximal dissipativity of $\hat{\Theta}^{-1}$. [^11] We point out that [24] uses a different notation. The operator $A_{\Theta}$ corresponds to $A_{\hat{\Theta}^{-1}}$ in the notation of [24].